The baggage truck A has a mass of 800 kg and is used to pull each of the 300-kg cars. If the tractive force F on the truck is = 480 N, determine:



a) The initial acceleration of the truck.


b) Forces in coupling B and C.


c) What is the acceleration of the truck if the coupling at C suddenly fails?


The car wheels are free to roll. Neglect the mass of the wheels.


answer asap.

1
Expert's answer
2021-11-22T17:02:02-0500

"F=ma"

"m =m_1+m_2 = 800+300=110000 kg"

Initial acceleration of the truck

"a =\frac{F}{m} =\frac{480}{110000}=0.00436 \frac{m}{s}"